Find the amount accumulated FV in the given annuity account. Hint [See Quick Example 1 and Example 1.] (Assume end-of-period deposits and compounding at the same intervals as deposits. Round your answer to the nearest cent.)

$1,500 is deposited quarterly for 20 years at 7% per year

FV = $
Answer #1

Future Value of annuity = A* ((1+rate)^n-1)/rate

= 1500*((1+7%/4)^(20*4)-1)/ (7%/4)

=1500*171.7938

= 257690.74